# Supermassive Binaries in Ultralight Dark Matter Solitons

### v2

@@ I Introduction
Ultralight dark matter (ULDM) is a dark matter candidate consisting of very light spin-0 particles [1, 2, 3, 4, 5, 6, 7, 8], motivated by the low mass scalars which arise in many string theoretic scenarios [9]. ULDM masses between $10^{-23}$ and $10^{-20}$ eV may resolve some apparent discrepancies between the predictions of cold dark matter [CDM] and observations on sub-galactic scalesmm thanks to the wavelike nature of ULDM, which smooths small-scale structure relative to CDM. By contrast, on scales much larger than the de Broglie wavelength of the underlying particle ULDM behaves identically to CDM, making it an attractive dark matter candidate.
@@ II Background
where $m$ is the mass of the ULDM particle, $r$ is the radial distance from the center, and $r_{c}$ is the core radius, at which the density drops to half its peak value. We take this as the initial soliton profile in our simulations.
@@ IV Results
A further effect not accounted for by simple semi-analytic models is the reshaping of the soliton as the binary approaches the center – the ULDM density profile becomes increasingly ‘pinched’ by the additional gravitational fields of the SMBH. At late times in our fiducial run, the soliton oscillates around a central density of $\sim 4\times 10^{7}\textrm{M}_{\odot}\textrm{pc}^{-3}$. This is five times higher than the initial central density, while the half-maximum radius falls to $\sim 1.1$pc, half the initial value. This is shown in Figure 9 and will increase the dynamical friction.
Figure 9: The ’pinching’ of the ULDM profile during the fiducial run.
@@
For a fixed soliton mass, the core radius scales as $r_{c}\propto m^{-2}$, where $m$ is the ULDM particle mass. Consequently, the central density scales as $m^{6}$. Via equation (11) we see that the dynamical friction is proportional to the density, so we expect a much more rapid decay at larger values of $m$. This will be offset to some extent by the velocity terms in the denominator but since $v\sim 1/\sqrt{r}$ for a point mass these will not rise as rapidly as the density term in the numerator. Figure 14 compares the evolution of black holes of mass $5\times 10^{7}\textrm{M}_{\odot}$ within a soliton of mass $10^{9}\textrm{M}_{\odot}$ for different values of $m$ confirming our expectations for decay rate, and we also see that the breathing mode is less pronounced at larger $m$.
@@ VII Conclusion and Discussion
Our simulations are simplified by starting with the SMBH binary moving inside a stationary soliton. Moreover, by omitting the wider halo we suppress the random walk of the soliton position driven by interactions with ULDM granules in the halo [29, 30], which will could reheat the SMBH binary. Consequently, it would be interesting (if computationally expensive) to simulate the full halo merger while accounting for the SMBH dynamics.

### v3

@@ I Introduction
Ultralight dark matter (ULDM) is a dark matter candidate consisting of very light spin-0 particles [1, 2, 3, 4, 5, 6, 7, 8], motivated by the low mass scalars which arise in many string theoretic scenarios [9]. ULDM masses between $10^{-23}$ and $10^{-20}$ eV may resolve some apparent discrepancies between the predictions of cold dark matter [CDM] and observations on sub-galactic scales thanks to the wavelike nature of ULDM, which smooths small-scale structure relative to CDM. By contrast, on scales much larger than the de Broglie wavelength of the underlying particle ULDM behaves identically to CDM, making it an attractive dark matter candidate.
@@ II Background
where $m$ is the mass of the ULDM particle, $r$ is the radial distance from the center, and $r_{c}$ is the core radius, at which the density drops to half its peak value. We take this as the initial soliton profile in our simulations. This configuration is somewhat idealised, as the recent merger between the parent galaxies will give a soliton that has not had time to relax to its ground state.
@@ IV Results
A further effect not accounted for by simple semi-analytic models is the reshaping of the soliton as the binary approaches the center – the ULDM density profile becomes increasingly ’pinched’ over the course of the simulation. This is due to the gravitational fields of the SMBH, which create a perturbation on the soliton that is not accounted in the initial stable profile. At late times in our fiducial run, the soliton oscillates around a central density of $\sim 4\times 10^{7}\textrm{M}_{\odot}\textrm{pc}^{-3}$. This is five times higher than the initial central density, while the half-maximum radius falls to $\sim 1.1$pc, half the initial value. This is shown in Figures 9 and 10, and will increase the dynamical friction.
Figure 9: The ‘pinching’ of the ULDM profile during the fiducial run.
Figure 10: Plots showing the ‘pinching’ of the soliton, at times 0, 0.016, and 0.162 Myr.
@@
For a fixed soliton mass, the core radius scales as $r_{c}\propto m^{-2}$, where $m$ is the ULDM particle mass. Consequently, the central density scales as $m^{6}$. Via equation (11) we see that the dynamical friction is proportional to the density, so we expect a much more rapid decay at larger values of $m$. This will be offset to some extent by the velocity terms in the denominator but since $v\sim 1/\sqrt{r}$ for a point mass these will not rise as rapidly as the density term in the numerator. Figure 15 compares the evolution of black holes of mass $5\times 10^{7}\textrm{M}_{\odot}$ within a soliton of mass $10^{9}\textrm{M}_{\odot}$ for different values of $m$ confirming our expectations for decay rate, and we also see that the breathing mode and pinching is less pronounced at larger $m$. The is consistent with the narrower profile and deeper central gravitational potential near the soliton centre at larger $m$, reducing the impact of the black hole potential.
@@ VII Conclusion and Discussion
Our simulations are simplified by starting with the SMBH binary moving inside a stationary soliton. Moreover, by omitting the wider halo we suppress the random walk of the soliton position driven by interactions with ULDM granules in the halo [29, 30], which will could reheat the SMBH binary. Consequently, it would be interesting (if computationally expensive) to simulate the full halo merger while accounting for the SMBH dynamics. Moreover, these simulations have focused on equal mass binaries but the single black hole case suggests that significantly unequal mass binaries may behave quite differently, potentially experiencing stone-skipping effects at large mass ratios. These possibilities all provide interesting directions for future investigation.